OpenStudy (anonymous):

what is the surface area of a sphere with a circumference of 30ft? round to the nearest tenth

OpenStudy (anonymous):

no need for a lengthy explination just a brief description and a answer please

OpenStudy (jdoe0001):

so we know that the sphere has a circumference of 30ft  keep in mind that \(\bf circumference= 2\pi r\qquad thus\implies 30=2\pi r\implies ?=r\)

OpenStudy (jdoe0001):

once you get the "r" adius you can get the \(\bf \textit{surface area}=4\pi r^2\)

OpenStudy (anonymous):

well my choices are 286 143 71 and 22

OpenStudy (jdoe0001):

so... what did you get for "r" though?

OpenStudy (anonymous):

I don't exactly understand how to get it

OpenStudy (jdoe0001):

\(\bf circumference= 2\pi r\qquad thus\implies 30=2\pi r\implies ?=r\) solve for "r" notice above

OpenStudy (anonymous):

could I just get an answer please? math dosnt suit me very well

OpenStudy (jdoe0001):

well... how come your teacher didn't provide the answer sheet along with the exercises though?

OpenStudy (anonymous):

probably forgot to print it off

OpenStudy (jdoe0001):

well... give him/her an email, ask him/her to send it over email or let him/her know that she forgot to print it, thus you didn't put it in

OpenStudy (jdoe0001):

and that should be fine

OpenStudy (anonymous):

lol I ask nicely for you to simply tell me a more than likely simple answer and your sarcastic

OpenStudy (jdoe0001):

sarcastic? where? I'm honest in saying that, ask him/her for it through email, he/she would understand

OpenStudy (anonymous):

why when you can just tell me now

OpenStudy (jdoe0001):

the other possibility is that, the teacher did not include it because you're expected to do it and thus understand it someone else doing it for you, so you can transcribe it over will only help you handwrite better, or type for that matter but will defeat the idea behind the exercise

OpenStudy (anonymous):

forget it your wasting my time if your not going to tell me then just stop bothering me